# Game Theory in Cosmology

### v1

@@ 5 Law of Universal Arbitrage Equilibrium
The Law of Universal Arbitrage Equilibrium: The late-time acceleration follows from the teleodynamic tendency of the Universe to evolve toward a state where the expansion no longer produces significant changes in the horizon-accessible microstate count. As the growth of these microstates stabilizes, the dynamics naturally approach an asymptotic accelerated phase, and dark energy is the macroscopic signature of this stabilization process.
@@ 6 Fates of a Teleodynamic Universe
The teleodynamic Friedmann system is first written in its bias–functional form, where the effective energy density and pressure are given by
@@
$\rho_{\rm TD}(t)\to\rho_{\rm TD,\infty}=\alpha\,\overline{\Phi}_{\infty},\qquad w_{\rm TD}\to-1$ (113)
@@
$\overline{\Phi}(t)\to 0,\qquad\varphi(x,t)\to 0$ (116)
@@
$\overline{\Phi}(t)<0,\qquad\rho_{\rm TD}(t)=\alpha\,\overline{\Phi}(t)<0$ (118)
@@
$\overline{\Phi}(t):\quad\text{growth}\;\longrightarrow\;\text{decay}$ (119)
@@
In passing, we should also now mention what the possible fates for the universe are for the bias functional in the form given in (60). The first natural fate for this ansatz is a de Sitter–like future, since although the functional contains a term that grows as $a^{n}$ at late times, the physical meaning of the parameter $\epsilon$ is the ability of the structure to continue generating new memory. In this regime, the homogeneous bias asymptotes to a constant, $\bar{\Phi}(t)\to\bar{\Phi}_{\infty}$, and the teleodynamic energy density settles to a fixed positive value. The teleodynamic equation of state then tends toward $w_{\rm TD}\to-1$, driving the Universe into a de Sitter-like attractor. The second class of asymptotic behaviors arises when the teleodynamic memory continues to build indefinitely, corresponding to a persistent, non-vanishing $\epsilon$ even at ultra late times. In this case, the term $a^{n}$ grows monotonically without bound, causing the homogeneous bias to increase without saturation. If the rate of memory accumulation is sufficiently strong, the energy density diverges in finite cosmic time, resulting in a Big Rip-type future in which the scale factor, Hubble rate, and teleodynamic bias all blow up.
7 What is Memory in the Teleological Perspective?
It is helpful to discuss the subtleties of what really constitutes “memory” in the teleodynamic perspective, given all the formulations we have discussed so far.
In the teleodynamic framework, memory is not a psychological or informational construct but a physical property of an evolving cosmic system. It represents the persistent imprint left on the large-scale gravitational field by the history of structure formation, matter flows, and tidal interactions across cosmic timescales. Because gravity in an expanding universe is long-range and non-ergodic, the distribution of matter does not “forget” its past configurations, which is to say that we cannot treat the physics here to be completely Markovian. Instead, every merger, filamentary collapse, void expansion, or tidal alignment leaves traces in the statistical and systemic organization of the cosmic web, and these persistent features accumulate into a global bias functional, which we here refer to as cosmological memory.
From a dynamical perspective, memory encodes how the Universe departs from the assumptions of perfect equilibrium that underlie standard statistical mechanics. This is exemplified by the fact that a collisionless, self-gravitating system never fully thermalizes, and its phase-space distribution maintains long-lived correlations that cannot be erased by expansion alone. These correlations are precisely what teleodynamics captures. As the Universe evolves, preferred organizational patterns survive Hubble timescales and feed back into both the background expansion and the growth of structure. In this sense, we say that memory arises as an emergent, non-local statistical field that quantifies how past gravitational organization biases present evolution.
Cosmologically speaking, this memory becomes especially important once structure formation becomes nonlinear, which is the case once structures have been forming for an extended period of time, as in the current age of the universe. Filaments, clusters, and voids act as reservoirs of persistent tidal information, storing the integrated effects of billions of years of gravitational interactions. This accumulated organization cannot be described by local density alone; instead, it manifests as a global functional of the cosmic web. Viewed this way, cosmological memory is not an additional field or exotic substance introduced by hand, but is instead the inevitable statistical and systemic outcome of a Universe that has evolved under long-range gravity, formed a hierarchy of structures, and retained correlations across cosmic timescales. Teleodynamics formalizes this intuition into a predictive framework as the accumulated memory of structure formation influences large-scale dynamics, shapes late-time acceleration, and modulates the growth of perturbations. In this sense, memory is the physical bridge connecting the microscopic history of gravitational organization to the macroscopic observables that define modern cosmology.

### v2

@@ 6 Law of Universal Arbitrage Equilibrium
The form of the pressure above also shows that because $\sigma_{\rm TD}\geq 0$ and $\dot{H}<0$ in an expanding Universe, the negative pressure driving acceleration emerges if and only if the accessible microstates at the horizon increase. This leads us to the Law of Universal Arbitrage Equilibrium: The late-time acceleration follows from the teleodynamic tendency of the Universe to evolve toward a state where the expansion no longer produces significant changes in the horizon-accessible microstate count. As the growth of these microstates stabilizes, the dynamics naturally approach an asymptotic accelerated phase, and dark energy is the macroscopic signature of this stabilization process.
7 Addressing the Coincidence Problem
A central puzzle in cosmology is the following question: Why do we happen to live “right now,” the one brief epoch when matter and dark energy densities are of the same order, even though they evolve very differently? This is known as the Coincidence Problem. In standard cosmology, this appears as an unexplained numerical accident between two fundamentally distinct components [108, 109, 110, 111, 112, 113, 114]. In the teleodynamic framework, however, we will now show that the horizon-thermodynamic formulation already shows that the effective negative pressure arises from non-equilibrium entropy production on the apparent horizon through the law of universal arbitrage equilibrium.
The starting point is the teleodynamic Raychaudhuri equation (6.5), together with the teleodynamic Friedmann relation (6.7) and the bias fluid closure for the effective energy density and pressure. These relations imply a dictionary between the bulk teleodynamic sector $(\rho_{\rm TD},p_{\rm TD})=\alpha(\overline{\Phi},\overline{\Pi}_{\rm TD})$ and the horizon quantities $(C_{\rm TD},\sigma_{\rm TD})$. In particular, it was shown above that in the late-time regime, one can write the teleodynamic pressure purely in terms of the horizon entropy production as (6.14) while the teleodynamic energy density takes the form
$\rho_{\rm TD}=\alpha\overline{\Phi}=-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\,\sigma_{\rm TD}$ (7.1)
We note here that at late times, nonrelativistic matter dominates the standard sector so that $p_{m}\simeq 0$, and the radiation terms $\rho_{r},p_{r}$ are negligible, so these expressions simplify accordingly.
To make the coincidence problem more explicit, it is useful to adopt the minimal entropy production law introduced earlier for late times, in which the non-equilibrium production rate is proportional to the degree of departure from de Sitter equilibrium as in (6.16). We note that (6.16) guarantees $\sigma_{\rm TD}\geq 0$ whenever $\dot{H}<0$ and ensures that entropy production vanishes smoothly as $\dot{H}\to 0$, so that the de Sitter attractor corresponds to a state of zero teleodynamic entropy production on the horizon. Substituting (6.16) into the last term of (7.1), one obtains
$\frac{H^{5}}{\dot{H}}\,\sigma_{\rm TD}=\frac{H^{5}}{\dot{H}}\left(2\xi\,\frac{-\dot{H}}{H^{3}}\right)=-2\xi H^{2}$ (7.2)
so that in the late-time regime with $p_{m}\simeq 0$ and negligible radiation the energy density (7.1) becomes
$\rho_{\rm TD}=-2M_{P}^{2}\dot{H}-\rho_{m}-\frac{M_{P}^{2}}{2C_{\rm TD}}\left(-2\xi H^{2}\right)=-2M_{P}^{2}\dot{H}-\rho_{m}+\frac{M_{P}^{2}\xi}{C_{\rm TD}}H^{2}$ (7.3)
It is now convenient to introduce the dimensionless combination
$\zeta\equiv\frac{\xi}{C_{\rm TD}}$ (7.4)
which really just encapsulates the ratio between the strength of non-equilibrium entropy production and the horizon microstate density. With this definition, the late-time teleodynamic energy density reads
$\rho_{\rm TD}=-2M_{P}^{2}\dot{H}-\rho_{m}+\zeta M_{P}^{2}H^{2}$ (7.5)
The bias-fluid Friedmann equation at late times can hence be written as
$3M_{P}^{2}H^{2}=\rho_{m}+\left(-2M_{P}^{2}\dot{H}-\rho_{m}+\zeta M_{P}^{2}H^{2}\right)$ (7.6)
so that the matter density cancels on the right-hand side and we obtain a closed evolution equation for $H$,
$3M_{P}^{2}H^{2}=-2M_{P}^{2}\dot{H}+\zeta M_{P}^{2}H^{2}$ (7.7)
Dividing through by $M_{P}^{2}$ and rearranging gives
$2\dot{H}=-(3-\zeta)H^{2}$ (7.8)
which shows that once the late-time teleodynamic entropy production law (6.16) is adopted, the Hubble rate evolves as if an effective constant equation of state were driving the universe. Equation (7.8) is a new result here in the sense that it encodes the late-time teleodynamic background dynamics entirely in terms of the single parameter $\zeta$ defined in (7.4).
To express the solution of (7.8) in terms of the scale factor, it is useful to use the relation $\dot{H}=(dH/da)aH$, and then substituting this into (7.8) leads to
$aH\frac{dH}{da}=-\frac{1}{2}(3-\zeta)H^{2}$ (7.9)
which can be written as
$\frac{d\ln H}{d\ln a}=-\frac{1}{2}(3-\zeta)$ (7.10)
Integrating (7.10) with respect to $\ln a$ gives us
$\ln H=-\frac{1}{2}(3-\zeta)\ln a+\text{const}$ (7.11)
so that the Hubble rate scales as
$H(a)=H_{0}\,a^{-\frac{1}{2}(3-\zeta)}$ (7.12)
where $H_{0}$ is the Hubble parameter at the present epoch, and we note that the scale factor gets defined by $a(t_{0})=1$, from here we can write
$H^{2}(a)=H_{0}^{2}\,a^{-(3-\zeta)}$ (7.13)
The total energy density is therefore
$\rho_{\rm tot}(a)=3M_{P}^{2}H^{2}(a)=3M_{P}^{2}H_{0}^{2}\,a^{-(3-\zeta)}\equiv\rho_{\rm tot,0}\,a^{-(3-\zeta)}$ (7.14)
with $\rho_{\rm tot,0}=3M_{P}^{2}H_{0}^{2}$. Also, the matter density will continue to obey its usual continuity equation, so in the nonrelativistic regime, one has
$\dot{\rho}_{m}+3H\rho_{m}=0\qquad\Rightarrow\qquad\rho_{m}(a)=\rho_{m0}\,a^{-3}$ (7.15)
where $\rho_{m0}$ is the present-day matter density, and since $\rho_{\rm tot}=\rho_{m}+\rho_{\rm TD}$, combining (7.14) and (7.15) gives the teleodynamic energy density as a function of the scale factor,
$\rho_{\rm TD}(a)=\rho_{\rm tot}(a)-\rho_{m}(a)=\rho_{\rm tot,0}a^{-(3-\zeta)}-\rho_{m0}a^{-3}$ (7.16)
Introducing the present-day balance
$\Omega_{m0}=\frac{\rho_{m0}}{\rho_{\rm tot,0}},\qquad\Omega_{{\rm TD},0}=\frac{\rho_{{\rm TD},0}}{\rho_{\rm tot,0}},\qquad\Omega_{m0}+\Omega_{{\rm TD},0}=1,$
one can write (7.16) as
$\rho_{\rm TD}(a)=\rho_{\rm tot,0}\left[a^{-(3-\zeta)}-\Omega_{m0}a^{-3}\right]$ (7.17)
The ratio of teleodynamic to matter energy densities is then
$\frac{\rho_{\rm TD}(a)}{\rho_{m}(a)}=\frac{\rho_{\rm tot,0}\left[a^{-(3-\zeta)}-\Omega_{m0}a^{-3}\right]}{\Omega_{m0}\rho_{\rm tot,0}a^{-3}}=\frac{1}{\Omega_{m0}}\,a^{\zeta}-1$ (7.18)
This simple expression is the key to the teleodynamic solution of the coincidence problem in the horizon entropy framework. It shows that once the late-time entropy production law (6.16) is adopted, the ratio $\rho_{\rm TD}/\rho_{m}$ grows as a power of the scale factor determined solely by the parameter $\zeta$ and the present-day matter fraction $\Omega_{m0}$.
With all this in mind, we finally come to the present epoch with $a=1$, and (7.18) reduces to
$\left.\frac{\rho_{\rm TD}}{\rho_{m}}\right|_{a=1}=\frac{1}{\Omega_{m0}}-1=\frac{\Omega_{{\rm TD},0}}{\Omega_{m0}}$ (7.19)
The epoch when teleodynamic and matter densities are equal is defined by the condition $\rho_{\rm TD}(a_{\rm eq})=\rho_{m}(a_{\rm eq})$, which, using (7.18) implies
$\frac{1}{\Omega_{m0}}\,a_{\rm eq}^{\zeta}-1=1$ (7.20)
$a_{\rm eq}^{\zeta}=2\Omega_{m0},\qquad\Rightarrow\qquad a_{\rm eq}=\left(2\Omega_{m0}\right)^{1/\zeta}.$ (7.21)
For typical observational values such as $\Omega_{m0}\simeq 0.3$, one has $2\Omega_{m0}\simeq 0.6$, so that for $\zeta$ of order unity, the equality scale factor $a_{\rm eq}$ lies close to the present epoch in logarithmic terms. For example, if $\zeta=1$ then $a_{\rm eq}\simeq 0.6$ and the corresponding redshift, using $1+z=1/a$, is $z_{\rm eq}\simeq 1/0.6-1\approx 0.67$, while for $\zeta=1.5$ one finds $a_{\rm eq}\simeq 0.7$ and $z_{\rm eq}\simeq 0.43$. These values are precisely in the range where observations indicate that cosmic acceleration turns on and where the ratio of dark energy to matter is of order unity. So we see that the coincidence problem is thus recast as the statement that $\zeta$ should be of order one, rather than as a requirement that two unrelated densities happen to cross at $z\sim\mathcal{O}(1)$.
But how justified are we to take $\zeta=\xi/C_{\rm TD}$ to be of order unity? Let us start by discussing both its numerator and denominator. The parameter $C_{\rm TD}$ is determined by the microstate density on the horizon and is given by $C_{\rm TD}=4\pi s_{\rm cell}/\ell_{\chi}^{2}$ in the microscopic construction of the teleodynamic horizon entropy, where $s_{\rm cell}$ is the entropy per correlation cell and $\ell_{\chi}$ is the coarse-grain length. For a teleodynamic universe that approximates Gibbons-Hawking normalization, one would have $C_{\rm TD}\sim 8\pi^{2}M_{P}^{2}$, but more generally, $C_{\rm TD}$ is set by the ratio between the horizon area and the correlation area, without invoking any extreme hierarchy. The parameter $\xi$ originates from the quadratic entropy production law $\sigma_{\rm TD}=\sum_{ab}\mathcal{L}_{ab}J_{a}J_{b}$ and encodes the effective strength of horizon-level Onsager coefficients in the coarse-grained theory. In the absence of fine tuning, one can expect the dimensionless combinations entering $\xi$ to be of order unity, just as in standard non equilibrium thermodynamics for many-body systems and so, the ratio $\zeta=\xi/C_{\rm TD}$ is naturally $\mathcal{O}(1)$ when measured in appropriate units, especially if cosmology selects a correlation scale $\ell_{\chi}$ comparable to the horizon scale or to a slowly varying function of it. From the observational side, the same parameter $\zeta$ controls the effective equation of state $w_{\rm eff}$ inferred from (7.8) so that matching the observed value $w_{\rm eff}\approx-1$ at late times requires $\zeta$ to be close to three, again an order unity value with no extreme tuning.
Putting these elements together, the horizon thermodynamic framework in cosmological teleodynamics addresses the coincidence problem in a conceptually economical way. The late time entropy production law (6.16) ties the growth of the teleodynamic energy density directly to the departure from de Sitter equilibrium as encoded in $\dot{H}/H^{3}$ and the resulting solution (7.18) shows that $\rho_{\rm TD}/\rho_{m}$ scales as $a^{\zeta}$ with $\zeta$ determined by the microscopic parameters of the teleodynamic horizon, and for any $\zeta$ of order unity the equality between teleodynamic and matter densities naturally occurs at a scale factor $a_{\rm eq}$ of order one. The epoch when $\rho_{\rm TD}\sim\rho_{m}$ is not an accidental numerical coincidence between two unrelated fluids, but the inevitable consequence of the teleodynamic scaling implied by horizon entropy production in a universe whose memory-bearing microphysics is characterized by $\mathcal{O}(1)$ parameters. So, the idea is that we live in an epoch when the non-equilibrium horizon entropy production induced by structure formation is at or near its maximum. It is only during this window that the teleodynamic sector naturally rises to a level comparable to the matter density. Before that, it is dynamically negligible, and after that, it simply behaves like a residual cosmological constant.
@@ 8 Fates of a Teleodynamic Universe
Finally, we consider the different fates of a teleodynamic universe. The teleodynamic Friedmann system is first written in its bias–functional form, where the effective energy density and pressure are given by
@@
$\rho_{\rm TD}(t)\to\rho_{\rm TD,\infty}=\alpha\,\overline{\Phi}_{\infty},\qquad w_{\rm TD}\to-1$ (8.1)
@@
$\overline{\Phi}(t)\to 0,\qquad\varphi(x,t)\to 0$ (8.4)
@@
$\overline{\Phi}(t)<0,\qquad\rho_{\rm TD}(t)=\alpha\,\overline{\Phi}(t)<0$ (8.6)
@@
$\overline{\Phi}(t):\quad\text{growth}\;\longrightarrow\;\text{decay}$ (8.7)
@@
In passing, we should also now mention what the possible fates for the universe are for the bias functional in the form given in (5.50). The first natural fate for this ansatz is a de Sitter–like future, since although the functional contains a term that grows as $a^{n}$ at late times, the physical meaning of the parameter $\epsilon$ is the ability of the structure to continue generating new memory. In this regime, the homogeneous bias asymptotes to a constant, $\bar{\Phi}(t)\to\bar{\Phi}_{\infty}$, and the teleodynamic energy density settles to a fixed positive value. The teleodynamic equation of state then tends toward $w_{\rm TD}\to-1$, driving the Universe into a de Sitter-like attractor. The second class of asymptotic behaviors arises when the teleodynamic memory continues to build indefinitely, corresponding to a persistent, non-vanishing $\epsilon$ even at ultra late times. In this case, the term $a^{n}$ grows monotonically without bound, causing the homogeneous bias to increase without saturation. If the rate of memory accumulation is sufficiently strong, the energy density diverges in finite cosmic time, resulting in a Big Rip-type future in which the scale factor, Hubble rate, and teleodynamic bias all blow up.
9 Subtleties of the Bias Functional
The central new object in the teleodynamic framework is the bias functional $\Phi$, introduced at the level of the effective action. So, the purpose of this short section is not to re-derive the field equations or consider the perturbation analysis (all of this has been dealt with in detail in the appendix), but rather to state clearly the minimal assumptions required for $\Phi$ to define a well-posed coarse-grained effective sector. This is to highlight the theoretical constraints that follow from covariance and cosmological symmetries, and to explain what is meant by "closure" at the level of linear perturbations without committing to a unique microscopic model. A separate forthcoming work of ours will be devoted to the bias functional itself, including systematic derivations of candidate microscopic and mesoscopic forms, controlled expansions, and physically motivated kernel constructions. Here, we restrict ourselves to the assumptions and constraints needed to interpret the present paper as a predictive framework rather than an arbitrary phenomenological reparameterization.
At the level of effective action, the foundational requirement is that $\Phi$ defines a diffeomorphism scalar functional so that $\sqrt{-g}\,\Phi$ is a generally covariant integrand, and this requirement is both necessary and highly constraining. It implies that $\Phi$ can depend only on the fields through scalar quantities, possibly including scalar combinations of coarse-grained matter variables and scalar environmental diagnostics, and, in particular, the dependence of $\Phi$ on the environment must itself be encoded in covariant constructions. Operationally, $\mathcal{E}$ may be considered to denote a finite collection of coarse-grained scalar descriptors of the macroscopic state, which are built from smoothed densities, expansion/shear scalars associated with coarse-grained flow, and tidal invariants. The key point is that environment dependence is not introduced by selecting a preferred frame by hand, but by specifying covariant scalar data that characterize the macroscopic gravitational state.
A second essential assumption is the existence of a controlled background-perturbation split. On an FLRW spacetime, any scalar functional evaluated on the cosmological solution admits a decomposition
$\Phi(t,\mathbf{x})=\bar{\Phi}(t)+\delta\Phi(t,\mathbf{x})$ (9.1)
where $\bar{\Phi}$ is the homogeneous mode and $\delta\Phi$ is the inhomogeneous perturbation with vanishing spatial mean in constant-time hypersurfaces. This is not a model assumption, but rather the standard cosmological statement that FLRW symmetries isolate the homogeneous sector at the background level. What is nontrivial is that the teleodynamic effective theory presumes that such a split remains meaningful after coarse-graining. Concretely, this means that the smoothing scale implicit in the construction of $\mathcal{E}$ is taken to be much smaller than the Hubble scale so that a well-defined homogeneous limit exists and residual inhomogeneities can be treated perturbatively. This "scale separation" is an implicit assumption behind any cosmological effective description, but it becomes particularly important when the functional is intended to encode environmental dependence. Note again that this field split has been discussed and rigorously justified in the appendix.
A third essential assumption is functional differentiability at the level required by the variational principle. This is important to note since the teleodynamic sector is defined through a standard metric variation of the effective action, $\Phi$ must be sufficiently regular such that its metric functional derivative exists in the sense required by the variational calculus at least on the cosmological backgrounds and perturbations considered. This excludes pathological constructions in which $\Phi$ depends on non-differentiable functionals of the metric or matter fields, and so it would mean in practice that this requirement is satisfied if $\Phi$ is built from differentiable scalar invariants and differentiable nonlocal smearings with sufficiently regular kernels. This assumption is mild, but it is also the precise mathematical statement of what it means for the teleodynamic sector to be an effective-field-theory contribution rather than an arbitrary modification of equations.
Another important aspect that we should be concerned with here concerns locality and the distinction between "nonlocal as a principle" and "nonlocal as a truncation". From the perspective of coarse-grained gravitational dynamics and our present framework, this allows $\Phi$ to be nonlocal in space and to carry finite memory in time, provided that this nonlocality is implemented covariantly. A convenient schematic representation is to think of $\Phi$ as depending on a set of scalar invariants evaluated locally and on covariant smearings of scalar sources over spacetime regions,
$\Phi(x)=\Phi\left(\mathcal{I}_{A}(x),\int d^{4}x^{\prime}\sqrt{-g(x^{\prime})}\,K_{A}(x,x^{\prime})\,\mathcal{J}_{A}(x^{\prime}),\int^{t}dt^{\prime}\,{\cal M}_{A}(t,t^{\prime})\,\mathcal{Q}_{A}(t^{\prime})\right)$ (9.2)
where $K_{A}$ and ${\cal M}_{A}$ are kernels and $\mathcal{I}_{A},\mathcal{J}_{A},\mathcal{Q}_{A}$ are scalar sources constructed from the macroscopic state. The point of writing (9.2) is not to choose specific kernels, but to make explicit the theoretical constraints, which are that the kernels must be such that covariance is preserved, the response is causal in the time direction, and the overall coarse-grained description remains stable and perturbative on the cosmological backgrounds of interest. The forthcoming dedicated work of ours will systematize the allowed kernel structures and their physical interpretations.
A very natural requirement also comes from causality and stability conditions. These ideas impose further general constraints that are worth explicitly stating because they delimit the admissible space $\Phi$ even before any data analysis. First, we can say that if $\Phi$ contains memory integrals, the corresponding kernels should be retarded so that $\Phi$ at time $t$ depends only on the macroscopic state at earlier times $t^{\prime}\leq t$. Second, one can note here that the effective teleodynamic sector must not introduce runaway behavior around the cosmological background, as at linear order, this is equivalent to requiring that the closure relations used to relate $\delta\Phi$ to the standard perturbation variables do not generate exponentially growing modes inconsistent with the existence of the observed near FLRW universe. Finally, it is also paramount that the teleodynamic contribution is intended as an effective coarse-grained deformation, and therefore, it should remain perturbative in the regime where the cosmological background solution exists. What this means, in terms of the theory’s structure, is that $\alpha\Phi$ and its perturbations should not invalidate the background expansion history assumed in the main analysis, except perhaps in explicitly identified controlled parameter regions. In a practical implementation, this would translate into sign and magnitude conditions on the leading linear response coefficients of the closure, but the underlying conceptual content is simply that $\Phi$ should define a stable, causal, and perturbative effective sector.
A final caveat one must keep in mind concerns closure and predictivity: the action principle fixes how $\Phi$ contributes once $\Phi$ is specified, but at the linearized level, one must still provide a relation between $\delta\Phi$ and the standard perturbation variables to close the system. Presently, we have treated this as a response-theory problem rather than as a microphysical model specification. The most general linear closure consistent with homogeneity and isotropy can be expressed as a causal response of $\delta\Phi$ to the matter and environment perturbations with retarded kernels. For the purposes of our present work, it is sufficient to state that such a closure exists and can be organized systematically in a controlled expansion, which can be done, for example, by adopting a quasi-instantaneous truncation in which the dominant response is local in time but may remain scale dependent in space. A simple yet very explicit closure example may then be taken as a finite range spatial response of Yukawa type,
$\delta\Phi(k,t)=\beta(t)\,\frac{1}{k^{2}+a^{2}(t)m^{2}(t)}\,\delta_{m}(k,t)$ (9.3)
where $\beta(t)$ would set the strength of the response and $m(t)$ sets an effective range. Note that equation (9.3) is not proposed as a fundamental form of $\Phi$, but as the simplest illustrative closure showing how predictivity can be achieved without specifying a unique microscopic bias functional. It also highlights what is constrained theoretically, which is that the response must be causal, it must remain perturbative, and its characteristic scales must admit a consistent separation from the background cosmology.
A common concern with effective frameworks is whether introducing a new structure leads to uncontrolled proliferation of free functions. In the teleodynamic formulation, this is not the case, as we see that at the background level, the teleodynamic sector introduces at most a single effective degree of freedom associated with the homogeneous mode $\bar{\Phi}(t)$, which in controlled limits reduces to a slowly varying vacuum-like contribution. At the level of linear perturbations, predictivity is retained by the fact that the response of $\delta\Phi$ to cosmological perturbations can be consistently truncated to a small number of linear response coefficients, analogous in spirit to the $\mu(k,a)$ and $\gamma(k,a)$ functions used in effective descriptions of modified gravity and dark energy. Importantly, these response functions are not introduced ad hoc here; rather, they arise from the action-level structure once a closure prescription is specified. Thus, teleodynamics does not introduce an arbitrary set of new unknowns, but rather reorganizes the description of late-time dynamics into a framework with a number of effective parameters comparable to existing approaches. This is done while still predicting a restricted pattern of signatures tied to the functional structure of $\Phi$.